# Partial Search in a Frozen Network is Enough to Find a Strong Lottery Ticket

### v2

@@ 3.1 SLT Existence in Frozen Networks
Here we provide a theoretical result indicating that an SLT capable of approximating a target network exists in a frozen network. This result is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof see Appendix A).
Lemma 3.1 (Subset-Sum Approximation in Randomly Locked Networks (Lemma A.2)).
@@
Then, by combining Lemma 3.1 and Lemma 2.2, we extend the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
Lemma 3.2 (Subset-Sum Approximation in Frozen Networks (Lemma A.3)).
@@
Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. [10] instead of Lemma 2.2, it follows that an SLT approximating a target network exists within a frozen network.
Theorem 3.3 (SLT Existence in Frozen Networks (Theorem A.4)).
@@
Let a target network $f_{T}$ with depth $L$ and a partially frozen source network $f_{S}$ with depth $L\!+\!1$ be given. Assume that the source network is randomly frozen with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer. Also assume that these networks have ReLU activation function and initialized by using a uniform distribution $U[-1,1]$. Then, except with exponentially small probability, a subnetwork $f_{SLT}$ exists in the frozen source network $f_{S}$ such that $f_{SLT}$ approximates the target network $f_{T}$ if the width of $f_{S}$ is sufficiently large.
Figure 4: Pruning and locking set the bounds of the search space, whose optimal position is investigated.
3.2 Partial SLT Search in Frozen Networks
This paper proposes a novel method for searching SLTs within frozen networks: partial SLT search. It only searches in part of the source network, as it has been pruned at initialization; and it only searches for part of the final SLT, as some of its parameters are predefined by locking.
Figure 4 illustrates how network freezing reduces the search space. The total amount of frozen parameters is determined by a global freezing ratio $F_{r}$, which is the sum of the respective pruning ratio $P_{r}$ and locking ratio $L_{r}$. Therefore, the SLT search space $S$ is formed by the $1\!-\!F_{r}$ non-frozen parameters of the original dense network. As explained previously, it is not possible to search for an SLT with a lower sparsity (denser) than the source network. Consequently, the pruning ratio sets the lower bound of the search space. On the other hand, it is not possible to search for an SLT that prunes more parameters than those available, so the locking ratio sets the upper bound of the search space. (Specifically, the upper bound is given by $1\!-\!L_{r}$.) Thus, these ratios allow to freely control the size and position of the search space over the SLT sparsity spectrum.
Frozen network construction:
We assume that the $l$-th layer of network with depth $L$ has weights $\bm{w}^{(l)}\in\mathbb{R}^{n^{(l)}}$. The number of frozen parameters is determined by the layer-wise freezing ratios $p_{f}^{(1)},...,p_{f}^{(L)}$, where each $p_{f}^{(l)}$ is defined as the sum of the layer-wise pruning ratio $p_{p}^{(l)}$ and the layer-wise locking ratio $p_{l}^{(l)}$, and $p_{f}^{(l)},p_{p}^{(l)},p_{l}^{(l)}\in[0,1]$.
To prune $p_{f}^{(l)}n^{(l)}$ parameters and lock $p_{l}^{(l)}n^{(l)}$ parameters, we generate two random masks, a pruning mask $\bm{m}^{(l)}_{p}\in\{0,1\}^{n^{(l)}}$ and a locking mask $\bm{m}^{(l)}_{l}\in\{0,1\}^{n^{(l)}}$, so that they satisfy $||1-\bm{m}^{(l)}_{p}||=p_{p}^{(l)}n^{(l)}$, $||\bm{m}^{(l)}_{l}||=p_{l}^{(l)}n^{(l)}$, and $(1-\bm{m}^{(l)}_{p})\cdot\bm{m}^{(l)}_{l}=0$ (i.e., we require pruning and locking to be implemented without overlap). The layer-wise weights frozen with these masks are calculated as
$\bm{w}^{(l)}_{f}$ $:=(\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})+\bm{m}_{l}^{(l)})\odot\bm{w}^{(l)}.$ (1)
These masks are fixed during training, and we search for SLTs only in the parts where $\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})$ is one.
Setting the layer-wise ratios:
Our method considers the following two existing strategies for determining the layer-wise pruning ratio from the desired global pruning ratio of the source network:
  • Edge Per Layer (EPL): Each layer’s pruning ratio is set so that they all have the same number of remaining weights [29, 10].
  • Erdős-Rényi-Kernel (ERK): The pruning ratio of layer $l$ is proportional to the scale $(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\times C_{out}^{(l)}\times k_{h}^{(l)}\times k_{w}^{(l)})$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively [7].
The same strategy is employed to determine the layer-wise pruning and freezing ratios from their respective global ratios, and then the locking ratios are calculated as the difference between the corresponding freezing and pruning ratios.
Freezing pattern encoding for model compression:
The freezing pattern can be encoded during inference as a ternary mask—a freezing mask—that indicates whether a parameter is pruned, locked, or part of the supermask. For example, encoding pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)\in\{-1,0,1\}^{n^{(l)}}$. Since this freezing mask is also random and fixed, it can be regenerated from its seed and ratios, similarly to the random weights. Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature [14, 26, 21]. Also, this memory size compression reduces costly off-chip memory access on specialized neural inference accelerators [14, 4], thus SLTs within frozen networks have the potential to perform even faster and more energy-efficient inference processing.
3.3 Optimal Positioning of the Search Space
Here we perform a preliminary investigation of the optimal positioning of the search space bounds for each given desired SLT sparsity by exploring different pruning:locking proportions with a fixed freezing ratio of the network. Figure 5 explores different configurations of a $80\%$ freezing ratio—a search space size of $20\%$ of the original dense source network—on a Conv6 network and compares the performance of the found SLTs on CIFAR-1010{10}10. As expected, the best-performing SLT of each prune:lock configuration is found at the center of the search space, which maximizes the number of candidate subnetworks. In other words, within a search space size $S$, the optimal SLT sparsity is $k=P_{r}+S/2$, where $P_{r}=\frac{\sum_{l}p^{(l)}_{p}n^{(l)}}{\sum_{k}n^{(k)}}$ is the global pruning ratio of the network. Conversely, for a given freezing ratio $F_{r}$ of the network and a desired SLT sparsity $k$, the optimal search space position is set by the pruning ratio $P_{r}=k-(1-F_{r})/2$ and the corresponding locking ratio $L_{r}=F_{r}-P_{r}$. In the cases where this would result in $P_{r}\!<\!0$ or $L_{r}\!<\!0$, we choose a best-effort approach that keeps the desired freezing ratio and sets the search space bounds to $P_{r}\!=\!0$ or $L_{r}\!=\!0$, respectively.
Additionally, Figure 5 compares the two strategies for setting ratios considered in Section 3.2, showing that EPL outperforms ERK in all cases. Consequently, hereafter the proposed method sets the global ratios in order to position the search space center as close as possible to the desired SLT sparsity, and then sets the layer-wise ratios using EPL.
@@ 4 Experiments
This section demonstrates that freezing reduces the search space in a broad range of situations by evaluating it on image classification and graph node classification. We evaluate the impact of the freezing ratio on various network architectures, identifying three scenarios. Then, we explore trade-offs between accuracy, search space size, and model memory size for different network widths.
@@ 4.1 Experimental Settings
The proposed partial SLT search is evaluated on image classification using the CIFAR-1010{10}10 [19] and ImageNet [31] datasets, and on node classification using the OGBN-Arxiv [16] dataset. CIFAR-1010{10}10 and ImageNet train data are split into training and validation sets with a 4:1 ratio, while for OGBN-Arxiv we use the default set split. We test the models with the best validation accuracy and report the mean of three experiment repetitions for CIFAR-1010{10}10 and OGBN-Arxiv, and the result of one experiment for ImageNet. Standard deviation of experiments conducted more than once is plotted as error bars.
@@
For image classification we employ the VGG-like Conv6 [32, 30], ResNet [13], and Wide ResNet [40] architectures, and for graph node classification the 4-layer modified GIN [37] architecture in Huang et al. [17], all implemented with no learned biases. ResNet and GIN use non-affine Batch Normalization [18], while Conv6 has no normalization layers. Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution [12].
@@
SLTs are searched using an extension of Edge-Popup [30] that enforces the desired sparsity globally instead of per-layer [33]. On CIFAR-1010{10}10, scores are optimized for $100$ epochs using stochastic gradient descent with momentum $0.9$, batch size $128$, weight decay $0.0001$, and initial learning rates of $0.01$ and $0.1$ for Conv6 and ResNet-1818{18}18, respectively. On ImageNet, scores are optimized by the same setting as ResNet-1818{18}18 on CIFAR-1010{10}10, but a $256$ batch size. On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW [23] with weight decay $0.0001$ and initial learning rate $0.01$. All experiments use cosine learning rate decay [22]. These can be adequately verified with two NVIDIA H100 SXM5 94GB GPUs.
4.2 Reducing the Search Space with Freezing
@@
An experiment is first conducted to investigate the effect of the freezing ratio of the network on SLT accuracy. We identify three scenarios, represented in Figure 6.
@@
In the cases where the optimal SLT sparsity is found at intermediate sparsity—e.g., around $50\%$ for Conv6 in LABEL:fig:sparsity_acc_conv6—pruning and locking can be applied with equally high ratios. Results show that the search space can be reduced by $40\%$ with small impact on accuracy, and by $70\%$ with still a moderate accuracy drop of $5$ points.
@@
Applying the much larger ResNet-1818{18}18 to the same task results in much stronger overparametrization. Therefore, optimal SLTs are found in the higher sparsity range, as revealed by LABEL:fig:sparsity_acc_resnet18, benefiting of much higher pruning than locking. Even though $90\%$ of the search space is reduced in this scenario, we can find $90\%$ sparse SLTs with $88.1\%$ accuracy.
@@
As an example of the scenario where optimal SLT sparsities are found in the denser range, benefiting from a higher locking ratio, we evaluate GIN in LABEL:fig:sparsity_acc_gin. Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the search space our method finds similarly performing SLTs of $69.8\%$ accuracy with $40\%$ sparsity.
@@
Interestingly, with low SLT sparsity (e.g., $10\%$ sparsity) in all three scenarios, despite the reduced search space, SLTs within frozen networks achieve higher accuracy than SLTs within dense networks. These results suggest that parameter freezing at an appropriate ratio has the effect of avoiding the inclusion of low-grade local optimal solutions in the search space. It is known that Edge-Popup struggles to find the SLTs of the global optimal solution [9]. While searching for SLTs in a dense network leads to convergence to a local optimal solution, a moderate reduction of the search space may reduce the number of less accurate local optimal solutions and facilitate convergence to a more accurate local optimal solution within the reduced search space. In other words, we conjecture that if the network is properly frozen, the local optimal solution for the reduced search space is close to the global optimal solution of the entire search space.
4.3 Accuracy-to-Search Space Trade-Off
Section 3 sets as a requisite of the subset-sum proof that the source network has a sufficiently larger width than the target network. Figure 6 explores varying the width of the source network to analyze its impact on accuracy and search space size. SLT sparsity is fixed to that of the best performing SLT found in a dense source network in Figure 6: $50\%$ in Conv6, $90\%$ in ResNet-1818{18}18, and $20\%$ in GIN.
The results show in all cases that the found SLTs in wide frozen networks for a given sparsity achieve similar or higher performance with a similar or smaller search space size compared to the dense source network. That is, the proposed method improves the accuracy-to-search space size trade-off.
As shown in LABEL:fig:search_space_acc_conv6 and LABEL:fig:search_space_acc_gnn, in the scenarios that only allow for medium to low pruning, there is a limit to the number of parameters that can be pruned for a given SLT sparsity, and only aggressive pruning leads to less accurate SLTs. By additionally locking parameters, our method reduces the search space further and finds SLTs of higher accuracy than those within sparse or dense networks for a similar search space size. Pruning only improves the trade-off in the scenario that allows higher pruning than locking, as shown in LABEL:fig:search_space_acc_resnet18, while our method improves the accuracy-to-search space trade-off in all scenarios.
Figure 7: Freezing achieves better accuracy-to-search space trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks. SLT sparsity set according to highest accuracy in Figure 6 with dense source networks.
@@ 4.3 Accuracy-to-Search Space Trade-Off
Figure 8: Compared to sparse ($\bullet$) or dense ($\bullet$) source networks (top-left is better), freezing achieves better accuracy-to-memory size trade-off.
4.4 Accuracy to Model Memory Size Trade-Off
@@
The freezing mask compression scheme proposed in Section 3.2 allows to reduce the memory model size during inference by regenerating both the random weights and the freezing mask with a random number generator (RNG). Here we consider this compression and investigate the accuracy-to-model size trade-off offered by the SLTs found by the proposed method. Furthermore, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization [30], which can be exploited for reduced computational cost in neural engines [14]. For SKC we scale weights by $1/\sqrt{1-k_{l}}$, where $k_{l}$ is the sparsity of each layer, as proposed by Ramanujan et al. [30].
@@
As illustrated in Figure 8, compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy-to-model size trade-off in all scenarios. In other words, partial SLT search not only reduces the search space during training, but also reduces the model size during inference, further improving performance over conventional SLT search methods.

### v3

@@ 3.3 SLT Existence in Frozen Networks
One question comes to mind here: does SLT exist that approximates a given target network, even if the parameters are randomly frozen (i.e., pruned or locked) in advance? It has been shown by Gadhikar et al. (2023) that an SLT that approximates a given target network exists in the pre-pruned source network if the source network is sufficiently wider and deeper than the target, but it is not known whether such an SLT exists in the situation of a frozen source network. Here, we provide a rough theoretical result indicating that an SLT capable of approximating a target network exists in a frozen network with a sufficiently large width. This result is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof, see Appendix A).
Lemma 3.1 (Subset-Sum Approximation in Randomly Locked Networks).
@@
Then, by using the same proof procedure as Lemma 3.1, we extend the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
Lemma 3.2 (Subset-Sum Approximation in Frozen Networks).
@@
Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. (2023) instead of Lemma 2.2, it follows that an SLT approximating a target network exists within a frozen network.
Theorem 3.1 (SLT Existence in Frozen Networks).
@@
Let a target network $f_{\textrm{T}}$ with depth $L$ and a partially frozen source network $f_{\textrm{S}}$ with depth $L\!+\!1$ be given. Assume that the source network is randomly frozen with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer, except for 00-th and $1$-st layers. Also assume that these networks use the ReLU activation function and are initialized with a uniform distribution $U[-1,1]$. Then, except with exponentially small probability, a subnetwork $f_{\textrm{SLT}}$ exists in the frozen source network $f_{\textrm{S}}$ such that $f_{\textrm{SLT}}$ approximates the target network $f_{\textrm{T}}$ if the width of $f_{\textrm{S}}$ is sufficiently large.
Note that this theorem is intended to provide an existential guarantee for the question raised at the beginning of this subsection, whether the freezing operation hurts the existence of SLTs that can approximate an arbitrary network; thus, we do not focus on the sharpness of the probability bound to keep the results simple. Nevertheless, proving sharper bounds may lead to capturing the effect of the freezing operation, which is left for future work.
@@ 4 Experiments
This section demonstrates that freezing reduces the SLT memory size in a broad range of situations by evaluating it on image classification and graph node classification. We evaluate the impact of the freezing ratio on various network architectures, identifying three scenarios. Then, we explore trade-offs between accuracy and model memory size for different network widths and architectures.
@@ 4.1 Experimental Settings
We evaluate the SLTs within frozen networks on image classification using the CIFAR-1010{10}10 (Krizhevsky, 2009) and ImageNet (Russakovsky et al., 2015) datasets, and on node classification using the OGBN-Arxiv (Hu et al., 2020) dataset. CIFAR-1010{10}10 and ImageNet train data are split into training and validation sets with a 4:1 ratio, while for OGBN-Arxiv we use the default set split. We test the models with the best validation accuracy and report the mean of three experiment repetitions for CIFAR-1010{10}10 and OGBN-Arxiv, and the result of one experiment for ImageNet. The standard deviation of experiments conducted more than once is plotted as error bars.
@@
For image classification we employ the VGG-like Conv6 (Simonyan & Zisserman, 2014; Ramanujan et al., 2020), ResNet (He et al., 2016), and Wide ResNet (Zagoruyko & Komodakis, 2016) architectures, and for graph node classification the 4-layer modified GIN (Xu et al., 2019) architecture in Huang et al. (2022), all implemented with no learned biases. ResNet and GIN use non-affine Batch Normalization (Ioffe & Szegedy, 2015), while Conv6 has no normalization layers. Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution (He et al., 2015).
@@
SLTs are searched using an extension of Edge-Popup (Ramanujan et al., 2020) that enforces the desired SLT sparsity globally instead of per-layer (Sreenivasan et al., 2022). On CIFAR-1010{10}10, scores are optimized for $100$ epochs using stochastic gradient descent with momentum $0.9$, batch size $128$, weight decay $0.0001$, and initial learning rates of $0.01$ and $0.1$ for Conv6 and ResNet-1818{18}18, respectively. On ImageNet, scores are optimized by the same setting as ResNet-1818{18}18 on CIFAR-1010{10}10, but a $256$ batch size. On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW (Loshchilov & Hutter, 2019) with weight decay $0.0001$ and initial learning rate $0.01$. All experiments use cosine learning rate decay (Loshchilov & Hutter, 2017). These can be adequately verified with two NVIDIA H100 SXM5 94GB GPUs.
4.2 Varying Desired SLT Sparsity at Different Freezing Ratios
@@
This section investigates the effect of the freezing ratio of the network on desired SLT accuracy. We identify three scenarios, represented in Figure 5.
@@
In the cases where the optimal SLT sparsity is found at intermediate sparsity—e.g., around $50\%$ for Conv6 in LABEL:fig:sparsity_acc_conv6—pruning and locking can be applied with equally high ratios. Results show that the supermask memory size can be reduced by $40\%$ with a small impact on accuracy, and by $70\%$ with still a moderate accuracy drop of $5$ points.
@@
Applying the much larger ResNet-1818{18}18 to the same task results in much stronger overparametrization. Therefore, optimal SLTs are found in the higher sparsity range, as revealed by LABEL:fig:sparsity_acc_resnet18, benefiting from much higher pruning than locking. Even though $90\%$ of the memory size is reduced in this scenario, we can find $90\%$ sparse SLTs with $88.1\%$ accuracy.
@@
As an example of the scenario where optimal SLT sparsities are found in the denser range, benefiting from a higher locking ratio, we evaluate GIN in LABEL:fig:sparsity_acc_gin. Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the memory size, our method finds similarly performing SLTs of $69.8\%$ accuracy with $40\%$ sparsity.
@@
Interestingly, with low SLT sparsity (e.g., $10\%$ sparsity) in all three scenarios, despite the reduced parameters to be optimized, SLTs within frozen networks achieve higher accuracy than SLTs within dense networks. These results imply that parameter freezing at an appropriate ratio has the effect of avoiding the inclusion of low-grade local optimal solutions in the search space. While searching for SLTs in a dense network by Edge-Popup leads to convergence to a local optimal solution (Fischer & Burkholz, 2022), a moderate random parameter freezing may reduce the number of less accurate local optimal solutions and facilitate convergence to a more accurate local optimal solution within the reduced search space. In other words, we conjecture that if the network is properly frozen, the local optimal solution for the reduced search space is close to the global optimal solution of the entire search space.
@@
Figure 6: Compared to sparse ($\bullet$) or dense ($\bullet$) source networks, freezing achieves better accuracy-to-model memory size trade-off (top-left is better).
4.3 Accuracy-to-Model Memory Size Trade-Off
@@
The freezing mask compression scheme proposed in Section 3.1 allows to reduce the model memory size during inference by regenerating both the random weights and the freezing mask with random number generators. Here we consider this compression and investigate the accuracy-to-model size trade-off offered by the SLTs found in the frozen networks. Model size refers to the total memory size of model parameters that need to be stored. The random weights and the frozen parts of the supermask are excluded, since they can be regenerated from seed, whereas each non-frozen element of the supermask and each learned batchnorm parameter are counted as 1 and 32 bits, respectively. Furthermore, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization (Ramanujan et al., 2020), which can be exploited for reduced computational cost in neural engines (Hirose et al., 2022). For SKC, we scale weights by $1/\sqrt{1-k_{l}}$, where $k_{l}$ is the sparsity of each layer, as proposed by Ramanujan et al. (2020). Note that since we are interested in the accuracy of the SLTs compressed as much as possible obtained by each method, the pre-pruning ratio for the pruning-only case is determined as $(\textrm{target sparsity}-5)\%$ to compress the SLT as much as possible while allowing to search for SLTs. On the other hand, we evaluate freezing with a higher compression ratio than pruning only. This setting means we evaluate our method under more strict conditions than pruning-only from the compression ratio perspective.
@@
Figure 6 explores varying the width of the source network to analyze its impact on accuracy and model size. SLT sparsity is fixed to that of the best performing SLT found in a dense source network in Figure 5: $50\%$ in Conv6, $90\%$ in ResNet-1818{18}18, and $20\%$ in GIN. Compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy-to-model size trade-off in all scenarios.